# Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy

### v1

@@ Abstract
We present the Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy. It is being virtually organized by the Research Center for Medical Image Analysis and Artificial Intelligence (MIAAI), Department of Medicine, Danube Private University, Krems, Austria and Medical Imaging and Signal Analysis Hub (MISAHUB) in collaboration with the 9th International Conference on Computer Vision & Image Processing (CVIP 2024) being organized by the Indian Institute of Information Technology, Design and Manufacturing (IIITDM) Kancheepuram, Chennai, India. This document describes the overview of the challenge, its registration and rules, submission format, and the description of the utilized datasets.
@@
Video Capsule Endoscopy, Abnormality Classification, Biomedical Challenge, Medical Image Analysis
Challenge related questions: ask.misahub@gmail.comRegistration form: LinkChallenge hosting website: LinkChallenge GitHub repository: LinkTraining and validation dataset: LinkSubmission report Overleaf: Link
@@ I Overview of the Challenge
A significant rise in the burden of Gastro-Intestinal (GI) and liver diseases across the globe has been observed due to varying environmental factors brought on by industrialization, changes in nutrition, and increased use of antibiotics [1, 2, 3]. Different endoscopic techniques are utilized in the diagnosis, treatment, and management of these diseases. Video Capsule Endoscopy (VCE) is one such technique which allows direct visualization of the GI tract and travels with the help of a disposable capsule-shaped device [4]. The device comprises an optical dome, a battery, an illuminator, an imaging sensor, and a transmitter.
VCE is a non-invasive technique, which does not suffer from any sedation-related complications and has improved a physician’s ability to find different anomalies in the GI tract - especially small bowel related diseases such as Crohn’s disease, Celiac disease, and intestinal cancer, among others [4, 5]. It is often considered as an alternative to conventional endoscopy techniques [5, 6].
The true potential of VCE technology is hindered by the longer reading time of the video frames without compromising the quality of the report and cost of the capsule [7]. During 6-8 hours of VCE procedure, a video of the GI tract trajectory is recorded on a device attached to the patient’s belt which produces about 57,000-1,000,000 frames [8]. These video frames are analysed retrospectively by experienced gastroenterologists. Presently, an experienced gastroenterologist takes approximately 2–3 hours to inspect the captured VCE video through a frame-by-frame analysis [8]. This analysis is subject to human bias and high false-positive rates due to bubbles, debris, intestinal fluid, foreign objects, and chyme (food) among other factors negatively affecting the mucosal frames [9]. Further, the inadequate doctor-to-patient ratio across the globe delays this process. Other hardware related technological limitations include capsule retention, battery limitations, and bowel obstructions [6, 7, 8].
@@ I-A Significance of the Challenge
  • The aim of the challenge is to provide an opportunity for the development, testing, and evaluation of AI models for automatic classification of abnormalities captured in VCE video frames.
  • This challenge consists of distinct training, validation, and test datasets for training, internal validation, and external validation purposes.
  • It promotes the development of vendor-independent and generalized AI-based models for an automatic abnormality classification pipeline with 10 class labels, namely: angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, worms, and normal.
@@ I-B Prizes
  • E-certificate to top three winning teams.
  • Co-authorship in the challenge summary paper for top 4 teams.
@@ II-A Rules for participation
  • This challenge will be open to all students (B. Tech/ M. Tech/ Ph.D. of all branches), faculty members, researchers, clinicians, and industry professionals across the globe for free.
  • It will take place in full-virtual mode.
  • Participants will either register as a solo participant or can form a team.
  • The results will be submitted in a specific format over email to ask.misahub@gmail.com. The template for submission has been provided.
  • In the case of ties, the organizing committee may rank teams based on the method’s novelty and readability of codes. The organizing committee’s decision in this regard will be final.
@@ II-B Rules for Team Formation
  • A team can have a maximum of 4 participants.
  • Team members can be from the same or different organizations/ affiliations.
  • A participant can only be a part of a single team.
  • Only one member from the team must register for the challenge.
  • Each team can only have a single registration. Multiple registrations per team can lead to disqualification.
  • There are no limitations on the number of teams from the same organizations/ affiliations. (However, each participant can only be part of a single team.)
@@ II-C Rules for use of Training and Validation Dataset
  • Store the model, associated weights and files. Compute the predictions and evaluation metrics on the validation dataset (internal validation) as per the provided sample code on our github repository.
  • The participants are allowed to perform augmentations, resize the data for saving computational resources, and can use existing machine learning, meta-learning, ensembles, any pre-trained models, and deep learning models.
  • A sample code has been provided on our github how to load the dataset.
@@ II-D Rules for use of Testing Dataset
  • Download the dataset. It consists of the un-labelled test dataset.
@@ II-E Submission format
Each team is required to submit their results in an EMAIL with the following structure to ask.misahub@gmail.com.
@@
    • Team member names and affiliation in the BODY OF THE EMAIL.
@@
    • A link of the github repository in public mode in the BODY OF THE EMAIL.
    • A link of their report on any open preprint server of their choice (ArXiv, Authorea, BioRxiv, Figshare etc) in the BODY OF THE EMAIL.
    • Generated excel sheet (in xlsx format) as an attachment.
@@
    • Developed code for training, validation, and testing in .py / .mat etc in readable format with comments.
@@
    • Any utils or assets or config. or checkpoints.
@@ II-F Important Notes
  • The challenge dates and the time zones will be as per the Central European Summer Time (CEST; UTC+02:00, Austria). They will be counted as per end of the day 23:59.
  • The github repository should be PUBLIC. Repositories which require access will NOT be considered for evaluation.
@@
  • The participants are requested to STRICTLY follow the submission format.
  • The participants are NOT allowed to utilize any other dataset for training/ validation/ testing in this challenge.
  • The participants are strictly advised to use the overleaf template for developing their report. The zip file is also available at our github repository.
  • The participants are requested to acknowledge and cite the datasets in their reports and for any other research purposes. Use of the datasets for commercial products or any purpose other than research is STRICTLY NOT ALLOWED.
  • The participants are encouraged to expand their novel AI pipeline to full-length manuscripts for consideration in prestigious conferences like CVIP 2024 and journals.
  • All participating teams allow the organizing team to summarize their developed AI pipelines in the form of a challenge summary paper.
  • Any submitted excel sheet with missing data will be evaluated but given zero marks.
  • For the final test set result submission, both training and validation datasets may be used for training.
  • For the final test set result submission, only one submission is allowed. In case of multiple submissions, only the last one will be considered.
  • Our team has conducted a preliminary analysis utilizing a variety of machine learning models. The sample codes are available on our github repository. We will release a standardized benchmarking pipeline along with its performance results on the validation dataset on our github repository coinciding with the release of the test datasets. Participants are requested to review these benchmarks, compare their own results accordingly, and include this comparative analysis in their submissions.
II-G Criteria of judging a submission
  • All emails received before the result submission closing date will be considered for evaluation.
  • The github repository, report on any open-source pre-print server, and excel sheet received for all entries
    will be downloaded by the organizing team.
  • A common data file will be prepared to compare the evaluation metrics achieved for testing the dataset.
  • A semi-automated python script file will be used to achieve the best evaluation metrics received among all entries.
  • The following checklist will be used to select the top three winning teams:
    • Best evaluation metrics (especially mean AUC) on testing dataset.
    • In case of ties:
        Model uniqueness and reproducibility.
        Readability of the submitted report and github repository.
        Unique methods of handling class-imbalance problems in the datasets.
III Datasets
III-A Training, validation, and test dataset description
The training and validation dataset has been developed using three publicly available (SEE-AI project dataset [11], KID [12], and Kvasir-Capsule dataset [13]) and one private (AIIMS [8]) VCE datasets. The training and validation datasets consist of 37,607 and 16,132 VCE frames, respectively, mapped to 10 class labels namely angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, worms, and normal. Fig. 1 and Fig. 2 depict the class distribution of the training and validation as per the source datasets. The source datasets in the training and validation dataset have ONLY been extracted, cropped, resized, and split for the development of a large dataset in this challenge.
Figure 1: Class-wise distribution of the training and validation datasets. The classes represent all the abnormalities including angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, and worms.
@@ III-A Training, validation, and test dataset description
Figure 3: Pipeline for developing the training and validation datasets for this challenge.
TABLE I: No. of VCE frames with respect to their class labels and source dataset in the training and validation datasets.
@@
The testing dataset has been developed using a high-quality, medically annotated private dataset from the Department of Gastroenterology and HNU, All India Institute of Medical Sciences Delhi, India. It consists of 4,385 VCE frames which have been collected from more than 70 patient VCE video with abnormalities namely angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, and worms. The images are de-identified and do not contain any patient information. The data collection for this AI study was done according to Helsinki declarations. It was approved by the Department of Gastroenterology and HNU, All India Institute of Medical Sciences Delhi, India ethics committee (Ref. No.: IEC-666/05.08.2022). Further it is part of the objectives of the Core Research Grant Project (CRG/2022/001755). Upon completion of the challenge, the test set will be open-sourced with its ground truths.
@@ III-B Development of the Training and Validation Dataset
EndoSLAM, AI-KODA, and rhode island dataset were not considered as they do not contain abnormalities and focus on virtual capsule reconstruction, cleanliness assessment, and organ classification respectively. Annotated bleeding dataset and red lesion endoscopy dataset were not considered as they only contain red lesions. WCEBleedGen was not considered as it contains frames from KID, kvasir-capsule, annotated bleeding dataset and only focuses on bleeding classification, detection, and segmentation. AutoWCEBleedGen-Test has been developed by the co-authors of this challenge and hence was considered as a part of the test dataset. All the frames have been anonymized and renamed.
Finally, the KID, kvasir-capsule, and SEE-AI project were considered. Then the VCE frames and their abnormalities in each of these datasets were downloaded, and extracted based on the inclusion and exclusion criteria.
@@
A new class label of worms with 226 VCE frames has been introduced in the training and validation dataset. VCE frames containing 95 ulcer VCE frames have also been added. The ulcer and worm VCE frames have been collected from three and three patients, respectively, from the department of gastroenterology, HNU, All India Institute of Medical Sciences Delhi, India and were utilized in the study conducted by Goel et al. [8].
All the frames were cropped and resized to $224\times 224224 × 224$. The frames were merged as per their source and class label folders and checked for discrepancies. Checks for discrepancies included that all VCE frames should be:
@@
After a thorough check, the main folder was subjected to a random data split of 70-30% for the development of the training and validation dataset. Table I details the number of VCE frames with respect to their respective class labels in each of the datasets.
IV Conclusion

### v2

@@ Abstract
We present the Capsule Vision 2024 Challenge: Multi-Class Abnormality Classification for Video Capsule Endoscopy. It was virtually organized by the Research Center for Medical Image Analysis and Artificial Intelligence (MIAAI), Department of Medicine, Danube Private University, Krems, Austria, and Medical Imaging and Signal Analysis Hub (MISAHUB) in collaboration with the 9th International Conference on Computer Vision & Image Processing (CVIP 2024) being organized by the Indian Institute of Information Technology, Design and Manufacturing (IIITDM) Kancheepuram, Chennai, India. This document provides an overview of the challenge, including the registration process, rules, submission format, description of the datasets used, qualified team rankings, all team descriptions, and the benchmarking results reported by the organizers.
@@
Video Capsule Endoscopy, Abnormality Classification, Biomedical Challenge, Medical Image Analysis, Artificial Intelligence
Important key links of the challenge:Challenge related questions: ask.misahub@gmail.comChallenge hosting website: LinkChallenge GitHub repository: LinkChallenge ArXiv: LinkTraining and validation dataset: LinkTest dataset: LinkSanity Checker: LinkSubmission report Overleaf: Link
@@ I Overview of the Challenge
A significant rise in the burden of Gastro-Intestinal (GI) and liver diseases across the globe has been observed due to varying environmental factors brought on by industrialization, changes in nutrition, and increased use of antibiotics [1, 2, 3]. Different endoscopic techniques are utilized in the diagnosis, treatment, and management of these diseases. Video Capsule Endoscopy (VCE) is one such technique that allows direct visualization of the GI tract and travels with the help of a disposable capsule-shaped device [4]. The device comprises an optical dome, a battery, an illuminator, an imaging sensor, and a transmitter.
VCE is a non-invasive technique that does not suffer from any sedation-related complications and has improved a physician’s ability to find different anomalies in the GI tract, especially small bowel-related diseases such as Crohn’s disease, Celiac disease, and intestinal cancer, among others [4, 5]. It is often considered as an alternative to conventional endoscopy techniques [5, 6].
The true potential of VCE technology is hindered by the longer reading time of the video frames without compromising the quality of the report and cost of the capsule [7]. During 6-8 hours of VCE procedure, a video of the GI tract trajectory is recorded on a device attached to the patient’s belt, which produces about 57,000-1,000,000 frames [8]. These video frames are analysed retrospectively by experienced gastroenterologists. Presently, an experienced gastroenterologist takes approximately 2–3 hours to inspect the captured VCE video through a frame-by-frame analysis [8]. This analysis is subject to human bias and high false-positive rates due to bubbles, debris, intestinal fluid, foreign objects, and chyme (food), among other factors negatively affecting the mucosal frames [9]. Further, the inadequate doctor-to-patient ratio across the globe delays this process. Other hardware-related technological limitations include capsule retention, battery limitations, and bowel obstructions [6, 7, 8].
@@ I-A Significance of the Challenge
  • The aim of the challenge was to provide an opportunity for the development, testing, and evaluation of AI models for automatic classification of abnormalities captured in VCE video frames.
  • This challenge consisted of distinct training, validation, and test datasets for training, internal validation, and external validation purposes.
  • It promoted the development of vendor-independent and generalized AI-based models for an automatic abnormality classification pipeline with 10 class labels, namely: angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, worms, and normal.
@@ I-B Prizes
  • E-certificate to all the ranked teams.
  • Co-authorship in the challenge summary paper for the top 4 teams.
@@ II-A Rules for participation
  • This challenge was open to all students (B. Tech/M.Tech/Ph.D. of all branches), faculty members, researchers, clinicians, and industry professionals across the globe for free.
  • It took place in full-virtual mode.
  • Participants were either registered as a solo participant or could form a team.
  • The results were asked to be submitted in a specific format over email to ask.misahub@gmail.com. The template for submission was provided.
@@ II-B Rules for Team Formation
  • A team could have a maximum of 4 participants.
  • Team members could be from the same or different organizations or affiliations.
  • A participant could only be a part of a single team.
  • Only one member of the team was allowed to register for the challenge.
  • Each team could only have a single registration. Multiple registrations per team led to disqualification.
  • There were no limitations on the number of teams from the same organizations or affiliations. (However, each participant could only be part of a single team.)
@@ II-C Rules for use of Training and Validation Dataset
  • Store the model, associated weights, and files. Compute the predictions and evaluation metrics on the validation dataset (internal validation) as per the provided sample code on our github repository.
  • The participants were allowed to perform augmentations, resize the data for saving computational resources, and could use existing machine learning, meta-learning, ensembles, any pre-trained models, and deep learning models.
  • A sample code was provided on our github how to load the dataset.
@@ II-D Rules for use of Testing Dataset
  • Download the dataset. It consisted of the unlabeled test dataset.
@@ II-E Submission format
Each team was required to submit their results in an email with the following structure to ask.misahub@gmail.com.
@@
    • Team member names and affiliations in the BODY OF THE EMAIL.
@@
    • A link to the github repository in public mode in the BODY OF THE EMAIL.
    • A link to their report on any open preprint server of their choice (ArXiv, Authorea, BioRxiv, Figshare, etc.) in the BODY OF THE EMAIL.
    • Generated Excel sheet of the predicted train, validation, and test dataset (in xlsx format) as an attachment.
@@
    • Developed code for training, validation, and testing in .py or .mat, etc. in readable format with comments.
@@
    • Any utils, assets, or config. or checkpoints.
@@ II-F Important Notes
  • The challenge dates and the time zones were considered to be as per the Central European Summer Time (CEST; UTC+02:00, Austria). They were counted as per end of the day 23:59.
  • The github repository should be public. Repositories that require access were NOT considered for evaluation.
@@
  • The participants were requested to STRICTLY follow the submission format.
  • The participants were NOT allowed to utilize any other dataset for training, validation, or testing in this challenge.
  • The participants were strictly advised to use the overleaf template for developing their report. The zip file is also available at our github repository.
  • The participants were requested to acknowledge and cite the datasets in their reports and for any other research purposes. Use of the datasets for commercial products or any purpose other than research is STRICTLY NOT ALLOWED.
  • All participating teams have allowed the organizing team to summarize their developed AI pipelines in the form of a challenge summary paper.
  • Any submitted Excel sheet with missing data was evaluated but given zero marks.
  • For the final test set result submission, both training and validation datasets were allowed for training.
  • For the final test set result submission, only one submission was allowed. In the case of multiple submissions, only the last one was considered.
  • The organizing team had conducted a preliminary analysis utilizing a variety of machine learning models. The sample codes were made available on our github repository.
  • All teams were asked to verify their generated excel sheets using the sanity checker.
Figure 1: Class-wise distribution of the training and validation datasets. The classes represent all the abnormalities, including angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyps, ulcers, and worms.
@@
Figure 3: Class-wise distribution of the testing dataset.
Figure 4: Pipeline for developing the training and validation datasets for this challenge. A similar pipeline was also utilized for development of the testing dataset.
TABLE I: No. of VCE frames with respect to their class labels and source dataset in the training, validation, and test datasets.
@@
Testing [12] AIIMS 37 334 49 17 48 155 3526 11 181 27
II-G Criteria of judging a submission
  • All emails received before the result submission closing date were considered for evaluation.
  • The github repository, report on any open-source pre-print server, and excel sheets received for all entries
    were downloaded by the organizing team.
  • A common data file was prepared to compare the evaluation metrics achieved for training, validation, testing the dataset. It is now available on our github.
  • A semi-automated Python script file was used to achieve the best evaluation metrics received among all entries and is available on our github.
  • The following checklist was used to select the top three winning teams:
    • Combined metric on testing dataset. The combined metric was calculated by averaging the mean AUC and balanced accuracy for each team, separately for each dataset. It yielded a value between 0 and 1, with higher values indicating better performance, reflecting a balance between how well the model distinguishes each class (mean AUC) and the accuracy across classes (balanced accuracy).
III Datasets
III-A Training, validation, and test dataset description
The training and validation dataset were developed using three publicly available (SEE-AI project dataset [13], KID [14], and Kvasir-Capsule dataset [15]) and one private (AIIMS [8]) VCE dataset. The training and validation datasets consisted of 37,607 and 16,132 VCE frames, respectively, mapped to 10 class labels, namely angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, worms, and normal. Fig. 1 and Fig. 2 depict the class distribution of the training and validation as per the source datasets. The source datasets in the training and validation datasets were ONLY extracted, cropped, resized, and split for the development of a large dataset in this challenge.
The testing dataset was developed using a high-quality, medically annotated private dataset from the Department of Gastroenterology and HNU, All India Institute of Medical Sciences, Delhi, India. It consisted of 4,385 VCE frames, which have been collected from more than 70 patient VCE videos with abnormalities, namely angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, and worms (see Fig. 3). The images were de-identified and does not contain any patient information. The data collection for this AI study was done according to Helsinki declarations. It was approved by the Department of Gastroenterology and HNU, All India Institute of Medical Sciences Delhi, India ethics committee (Ref. No.: IEC-666/05.08.2022). Further, it is part of the objectives of the Core Research Grant Project (CRG/2022/001755). The ground truth labels of the test dataset are now released and updated on Figshare.
@@ III-B Development of the Training and Validation Dataset
EndoSLAM, AI-KODA, and rhode island datasets were not considered as they do not contain abnormalities and focus on virtual capsule reconstruction, cleanliness assessment, and organ classification, respectively. Annotated bleeding dataset and red lesion endoscopy dataset were not considered as they only contain red lesions. WCEBleedGen was not considered as it contains frames from KID, kvasir-capsule, annotated bleeding dataset and only focuses on bleeding classification, detection, and segmentation. AutoWCEBleedGen-Test has been developed by the co-authors of this challenge and hence was considered as a part of the test dataset. All the frames were anonymized and renamed.
Finally, the KID, kvasir-capsule, and SEE-AI projects were considered. Then the VCE frames and their abnormalities in each of these datasets were downloaded and extracted based on the inclusion and exclusion criteria.
@@
TABLE II: Results sorted in descending order based on the combined metric calculated for the testing set by the organizing team.
Rank Team Name Mean AUC Balanced Accuracy Combined Metric
1 PuppyOps [22] 0.8570334 0.3573283 0.6071809
2 MedInfoLab IIT Hyderabad [23] 0.7736101 0.3719341 0.5727721
3 WueVision [24] 0.7625353 0.3710301 0.5667827
4 Llama_Mamba [25] 0.7632455 0.3366313 0.5499384
5 Seq2Cure [26] 0.7460799 0.3467811 0.5464305
6 Taaldhwaj [27] 0.7271482 0.3359341 0.5315412
7 Capsule Commandos [28] 0.7313885 0.3234582 0.5274234
8 eAI [29] 0.7486922 0.2759186 0.5123054
9 VCap [30] 0.7364247 0.2813258 0.5088752
10 DS & Chill 0.8197379 0.1918587 0.5057983
11 Layer Players [31] 0.6585702 0.3509284 0.5047493
12 Rookies [32] 0.7237849 0.2617322 0.4927586
13 Team_CSIR [33] 0.6968434 0.2465738 0.4717086
14 CapsuleNet [34] 0.6724393 0.2674035 0.4699214
15 Code Cortex [35] 0.6821873 0.2298121 0.4559997
16 ViFo Tech [36] 0.6166294 0.2826887 0.4496591
17 1b2w [37] 0.6426787 0.2363980 0.4395384
18 Machine Minds [38] 0.6980735 0.1786562 0.4383648
19 aiVengers 0.6432135 0.2270782 0.4351459
20 Optiminds 0.7113022 0.1541079 0.4327051
21 Pioneers [39] 0.7065916 0.1195959 0.4130937
22 Organic 0.6011660 0.2231773 0.4121716
23 STEM sisters [40] 0.5821777 0.1398035 0.3609906
24 DeepScope Innovators [41] 0.5632873 0.1083343 0.3358108
25 BotBotBot [42] 0.5212472 0.1354561 0.3283516
26 Deep_Learners [43] 0.5118418 0.1194708 0.3156563
27 EndoAI [44] 0.4974579 0.0766949 0.2870764
TABLE III: Benchmarking results calculated by the organizing team on the test dataset.
S. No. Model Name Mean AUC Balanced Accuracy Combined Metric
1 VGG19 0.5254758 0.1444602 0.3349680
2 Xception 0.5341090 0.1313365 0.3327228
3 ResNet50V2 0.5421599 0.1772507 0.3597053
4 MobileNetV2 0.5484505 0.1140326 0.3312415
5 InceptionV3 0.5249649 0.1284446 0.3267047
6 InceptionResNetV2 0.5232011 0.1469184 0.3350598
A new class label of worms with 226 VCE frames was introduced in the training and validation dataset. VCE frames containing 95 ulcer VCE frames were also added. The ulcer and worm VCE frames were collected from three different patients, respectively, from the department of gastroenterology, HNU, All India Institute of Medical Sciences, Delhi, India, and were also utilized in the study conducted by Goel et al. [8].
All the frames were cropped and resized to $224\times 224224 × 224$ pixel resolution. The frames were merged as per their source and class label folders and checked for discrepancies. Checks for discrepancies included that all VCE frames should be:
@@
After a thorough check, the main folder was subjected to a random data split of 70-30% for the development of the training and validation dataset. Table I details the number of VCE frames with respect to their respective class labels in each of the datasets. Similar standard pipeline was followed for the development of the test dataset.
IV Results
This challenge received about 150 team participations across the globe during the registration stage, out of which 35 teams finally submitted their results. After resolving several discrepancies in initial submissions and requesting corrected files, the organizing team finalized a total of 27 qualified teams for evaluation. Table II details the rankings calculated by the organizing team of 27 qualified entries based on their performance on the test dataset. Table III details the benchmarking results calculated by the organizing team on the test dataset. No data augmentation, nor fine-tuning of the the AI models was performed in the benchmarking experiments. We are providing a description of all the 35 received entries in the Annexure A and also release the results for the train, validation, and test datasets for all teams, and our benchmark experiments through Table IV and V. All supporting codes, received excel sheets, reports of all teams and result evaluation done by the organizing team is available on our github.
V Conclusion